OpenStudy (anonymous):

help please! find the interest. principal, $400 rate, 5% time, 1 year

OpenStudy (mertsj):

The formula is Interest = principal multiplied by rate multiplied by the time

OpenStudy (anonymous):

\[S.I=\frac{ P \times R \times T }{ 100? }\]

OpenStudy (anonymous):

S.I=simple interest P=principal R=rate T=time